The resistance of a platinum coil of platinum resistance thermometer at zero degrees celcius is 5.0 ohms and at hundred degrees celcius is 5.39 ohms. When the thermometer is put in a hot bath, the resistance of platinum wire is found to be 5.795 ohms. What is the temperature of the bath


1
Expert's answer
2021-03-08T08:25:25-0500
"\frac{R-R_1}{R_2-R_1}=\frac{T-T_1}{T_2-T_1}\\
\frac{5.795-5}{5.39-5}=\frac{T-0}{100-0}\\T=203.8\degree C"